# Spurious reconstruction from brain activity

### v3

@@ Results
Pairwise identification has been a standard metric for evaluating latent feature decoding [19] or reconstruction performance [39, 29, 38, 34, 36, 40, 41, 44, 45, 46, 57, 58]. However, our analysis revealed that even with difficulties in accurately identifying semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig 6F). Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
@@
Pairwise identification is calculated as the accuracy with which the predicted features (either the output of a translator or features extracted from the output of a generator) can correctly identify the corresponding true ones, in pairs consisting of a true sample and one of the remaining samples in the test set. If the candidate sample belongs to the same category as the true sample, the identification is expected to be difficult. Conversely, if the candidate sample belongs to a different category from the true one, identification becomes easier.
@@
Fig 10: Expected pairwise sample identification performance in categorically structured data. The $x$-axis represents the number of categories in the test set. The $y$-axis represents the pairwise sample identification accuracy. Different lines represent various levels of accuracy ($q$) in distinguishing samples from different categories. The chance level of $0.5$ is represented by the bottom line. Above the graph, hypothetical similarity matrices are shown to illustrate the categorical structures of the samples, where the color (white/black) indicates similarity between samples. The block diagonal structure reflects the categorical nature of the data. Samples within the same category are assumed to be indistinguishable, resulting in a pairwise identification accuracy of $0.5$ (chance level). Samples from different categories can be distinguished with a pairwise identification accuracy of $q$, where q varies between $0.5$ and $1$.
@@
In this model, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity. We model the situation mentioned above by setting the expected identification accuracy at the chance level (i.e., $0.5$) when a candidate sample belongs to the same category as the true one. Conversely, when the candidate sample belongs to a different category than the true sample, we set the expected identification accuracy to a parameter $q$, ranging from 0.5 to 1. This parameter $q$ reflects the ease of identification across categories. Assuming the number of test samples is sufficiently large, the pairwise identification accuracy $\mathrm{Acc}$ becomes
@@
(See Materials and Methods “Expected identification accuracy in imprecise reconstructions” for the derivation).
@@
Fig 10 illustrates the relationship between pairwise identification accuracy and the number of categories in the test set through line plots of expected values. Inset figures depict the underlying similarity structure of the test set. Notably, even when identification within categories fails completely and succeeds only between two categories, the pairwise identification accuracy can still reach a high value of up to $75$%. This finding underscores a significant limitation of the pairwise identification metric: evaluating reconstruction performance based solely on whether this metric is above the chance level may lead to misleading conclusions.
Preserved image information across hierarchical DNN layers
@@
The reconstruction of arbitrary visual images requires compositional latent features that can be effectively mapped into the image space. Hierarchical DNN features are often employed as such representations. However, do these features truly map to the image space? A common narrative suggests that the hierarchical processing discard pixel-level information through progressively expanding receptive fields. This view, however, is not entirely accurate.
@@
Fig 11: Image recovery check for hierarchical latent features. The images show the ability to recover visual inputs from different layers of the VGG19 network. For each of the three sample images, three different optimization methods were applied. The left column shows results from pixel optimization, which directly optimizes pixel values to minimize the loss between the generated image features and the target image features [65]. The middle column displays optimization with a weak image prior, simultaneously optimizing both the weight parameters of the image prior models and their latent features [68]. The right column presents optimization with a pre-trained image prior, which optimizes the latent features of a parameter-fixed image generator model to minimize the loss between the output image features and the target image features [69]. Rows correspond to different layers of the VGG19 network, progressing from earlier layers at the top to later layers at the bottom. This progression illustrates how image information is preserved or lost at different stages of the network. Reasonable image recovery is possible even from relatively high-level layers of the VGG19 network, challenging the notion that higher layers in deep neural networks discard image-level visual information.
@@
For example, the latent features of auto-encoder models [63, 51, 64] can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the model’s output is trained to match the input. Furthermore, Mahendran and Vedaldi [65] showed that input images can be reconstructed with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task. It has also been argued in the neuroscience field that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant [66, 67]. These results challenge the notion that higher-level layers in DNNs discard all pixel-level information.
@@
To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Fig 11; see also Fig 7). Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (See Materials and Methods “Recovery check of a single layer by iCNN”). We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layer out of the total 19th layers). Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers. By utilizing a weak image prior [68], which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer. When using an image generator that has learned natural image information [29, 69], input images can be recovered even from the 15th layer.
@@
Conversely, utilizing high-level image features, such as the output of DNNs, or features from other modalities, such as text annotations, is not a rational choice for visual image reconstruction tasks. These latent features make it challenging to recover the corresponding input image (Figure 7B) and are insufficient as surrogate representations of perceived visual images. Although recent text-to-image models and predicted text features can easily generate images, the outputs should not be interpreted as reconstruction results. Instead, it is more appropriate to view them as visualizations of decoded semantic information. Such visualizations are valuable for illustrating purposes. However, it is crucial to recognize the significant distinction between semantic visualization and reconstruction.
How are we fooled by hallucinations of generative AIs?
Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and realistic images from text input [23] or generating text of a quality indistinguishable from human-written content [24]. However, due to the complex internal structure of these models and the vast amounts of data they are trained on, we are often fooled by the outputs of generative AIs. For instance, when searching for an unfamiliar topic using a large language model (LLM) in our daily lives, we may not realize that the model is creating false concepts. This is likely a result of generative AIs being trained on a large amount of data and producing highly coherent and contextual responses. We may also believe these models are unbiased and can represent all possible data points, even though they inherently contain biases from their training data and developers [70]. As we have observed, the generative AI-based reconstruction methods exhibit realistic appearances but poor generalizability (Fig 2). Could similar issues occur in visual image reconstruction studies as well?
@@
Traditionally, we have held two beliefs: (1) generating realistic images from brain activity is challenging, and (2) if the reconstruction pipeline effectively captures the brain’s representation under natural image perception, the model’s output should also appear realistic. Based on these beliefs, realistic reconstruction is often considered as an indication of accurately reflecting the actual visual experience.
Fig 12: The illustration of how realistic appearances can be misleading in visual reconstruction. On the left, Venn diagrams depict the relationship between two events: $R$, where the output has realistic appearances, and $T$, where the output truthfully reflects the visual experience. $T$ and a small dashed circle of $R$ depict a common pre-generative AI heuristic: the assumption that realistic appearance implies truthful reconstruction. Dashed arrows illustrate the effect of generative AI models (e.g., diffusion models) on this relationship. These models can potentially increase the overlap between $T$ and $R$, but may also expand areas of $R$ unrelated to $T$. On the right, a probabilistic interpretation of the relationship is presented. The equation represents the probability that the generator’s output truthfully reflects visual experience ($T$), given that the generator’s output has a realistic appearance ($R$). This conditional probability is derived from the areas in the Venn diagram. The expansion of $R$ without a proportional increase in $T$ can lead to a decrease in $\Pr[T\mid R]$, emphasizing the need for careful evaluation of reconstruction results beyond just assessing their realistic appearance.
@@
However, recent developments in generative AIs, such as diffusion models, have made it easy to produce convincing, realistic outputs, subverting the first assumption, i.e., $\Pr[R]\gg\Pr[\bar{R}]$. Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear realistic. Rather, as shown in Fig 12, the probability $\Pr[T|R]$ may become smaller as the generative AIs produce more convincing outputs (see also Fig 7). This perspective emphasizes the need for careful evaluation of reconstruction performance, considering the possibility of hallucinations by generators. While pursuing realistic reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
Discussion
@@
While our study critically examined the limitations of recent text-guided, diffusion-based methods for visual image reconstruction, it is important to acknowledge that these approaches provide new and potentially promising directions for brain decoding research. Although we emphasized the importance of achieving zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks, which, while not zero-shot, have provided insights into neural representations [1, 2]. Moreover, the visualization of semantic contents (e.g., the supplementary movies in Horikawa et al. [4]), can have significant utility in visual communication, even if it does not constitute zero-shot prediction. It is also worth noting that the individual components of these text-guided, diffusion-based methods are already being utilized in various brain decoding applications. For instance, text latent features derived from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity [71, 72, 73]. Furthermore, diffusion models are not limited to text-to-image generation; they can be trained to generate images only from visual latent features, as Cheng et al. [32] demonstrated in successful reconstruction of subjective experiences. By leveraging these components and exploring their potential synergies, researchers can continue to push the boundaries of brain decoding and visual image reconstruction, while keeping in mind the challenges and limitations highlighted in our study.
@@
The recent trend of collecting and sharing large-scale visual neural datasets, such as those by Hebart et al. [14] and Xu et al. [15], is a welcome development in the field of neuroscience. These datasets provide valuable resources for researchers to investigate brain function and advance our understanding of visual processing. The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli [12, 13]. The NSD has been widely utilized in various studies [74, 75, 76], demonstrating its value to the research community. While our results suggest that the semantic and visual diversity of the NSD stimuli may not be as high as initially thought, and there is substantial overlap between the training and test sets provided by the NSD authors, this does not diminish the overall importance and usefulness of the dataset. However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to consider the data split between training and test sets carefully. When aiming for generalizable predictions, such as in visual image reconstruction, researchers should verify whether significantly similar stimuli are included in both the training and test sets. If necessary, redesigning the training and test split can ensure that the models are tested on their ability to predict novel stimuli. Moreover, recent advancements in functional alignment and inter-site neural code conversion methods [77, 78, 79] hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience. These techniques allow researchers to align brain activity patterns across individuals and measurement sites even when stimuli are not shared across datasets. By leveraging these methods, researchers can pool data from various sources, increasing the sample size and diversity of the combined dataset, mitigating the limitations of individual datasets, and enhancing the development of generalizable and zero-shot prediction models.
@@
Investigating neural responses to natural stimuli is a highly valuable approach to understanding brain function and representation [80, 81]. As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli, especially in model training. However, we should not forget that we are also capable of perceiving non-natural stimuli like artificial images. We would like to emphasize that there are potential pitfalls when relying too heavily on evaluations based solely on natural stimuli. With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the text-guided reconstruction methods exploited the semantic and visual overlap between training and test sets. In the field of comparative and developmental psychology, researchers often prioritize using not natural but simple stimuli for better experimental controls and more precise inferences about infants’ cognitive abilities [82, 83]. Drawing inspiration from this approach, we argue that the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone. While natural stimuli are essential for ensuring ecological validity and understanding how the brain processes real-world information, it is equally important to assess the performance in controllable and transparent manners.
@@
We have observed that an inappropriate split between training and test stimulus sets can lead to spurious reconstruction, invalidating zero-shot predictions. It is crucial to recognize that these problems can arise in various research contexts. Several studies have used data where the stimuli shared the category information between the training and test sets, potentially compromising the validity of their results. Kavasidis et al. [84] collected a dataset of EEG signals recorded during natural image perception. Their visual stimuli consisted of $2,000$ images selected from $40$ object categories in ImageNet ($50$ images per category) and the test set contained the same categories that are included in the training set (See also [85, 15] for other issues with the data set). Denk et al. [57] attempted to develop music reconstruction methods from fMRI activity patterns. Their music stimuli consisted of $540$ music pieces selected from $10$ music genres, and the test set contained the same genres as the training set [86]. Orima et al. [87] attempted reconstructing perceived texture images from EEG signals. Their texture stimuli consisted of $191$ image patches extracted from $21$ natural textures, and they performed a reconstruction analysis in a leave-one-out manner. It should be carefully examined whether these studies may suffer from output dimension collapse, merely decoding the information from trained categories from brain activity.
@@
Experimental design for train-test stimuli setups requires careful consideration. Dado et al. [88] conducted an image reconstruction analysis from the multi-unit activity of a macaque using images generated from latent features of generative models. Importantly, all of the test stimuli were generated from the averaged latent features of the categories used in their training phase. While the authors addressed potential biases by redesigning the training and test split, researchers should exercise prudence when utilizing the dataset. Our inspection of the movie stimuli from Nishimoto et al. [89] revealed that many frames in the test movie stimuli were nearly identical to those in the training set (S5 Fig). This similarity likely results from temporally adjacent video frames being split between the training and test stimuli. While our preliminary analysis of their latent features (motion energy features) did not show unusual clustering (note also that the study presents a way of retrieving movie instances via a brain encoding model, rather than reconstruction in the current sense), caution is required when using the stimulus set to extract other types of features [90, 91]. Claims of predicting arbitrary instances or achieving zero-shot prediction warrant thorough scrutiny.
@@
There is a widespread practice of using test data for fine-tuning, which can also be a questionable procedure. Some studies have proposed methods that involve fine-tuning models using the entire test brain data, but not the test stimuli, following initial training with the training stimuli and brain data [39, 43]. While the absence of test labels (stimuli) during fine-tuning may help avoid obvious overfitting, these approaches still treat test brain data as previously observed information, rather than as a proxy for novel data potentially encountered in real-world situations. Consequently, this practice violates test data independence, making it difficult to evaluate the model’s actual generalization ability. Although such procedures may effectively improve performance on existing benchmarks or competitions, it is crucial to recognize that incorporating test data information during any stage of training can undermine the validity of neuroscientific claims and limit the real-world applicability of the methods.
@@
The issue of double dipping, which refers to the use of the same dataset for both data/variable selection and selective analysis (inference and prediction), has been widely recognized in neuroscience [92, 93]. In classification tasks, while selecting input variables using test data is problematic, using the same output labels (target variables) for training and test sets is not inherently flawed, given that the nature of classification assumes consistent categories across datasets. However, the challenges we address in this study, although related to double dipping, present distinct concerns. In evaluating zero-shot prediction performance, the mere similarity of test labels (features) can lead to overestimating model performance. It is crucial to distinguish between conventional double dipping and the current issues we identify, such as output dimension collapse in zero-shot prediction scenarios. These emerging challenges necessitate careful consideration not only of input data independence but also of structural similarities between training and test sets in the output space.

### v5

@@ 2.2.3 Caveat with evaluation by pairwise identification
Pairwise identification has been a standard metric for evaluating latent feature decoding (Horikawa and Kamitani, 2017) or reconstruction performance (Beliy et al., 2019; Shen et al., 2019a; Mozafari et al., 2020; Qiao et al., 2020; Ren et al., 2021; Gaziv et al., 2022; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023; Scotti et al., 2023; Denk et al., 2023; Koide-Majima et al., 2024). However, our analysis revealed that even with difficulties in accurately identifying specific semantic clusters the test samples belong to, the pairwise identification performance still surpassed its chance level (Fig. 6F). This result highlights a fundamental limitation of pairwise identification: its susceptibility to overestimation when the dataset or target features contain a strong high-level categorical structure. Here we critically examine this metric and demonstrate that significant results can be easily obtained when the target or predicted features exhibit certain structures.
@@
Pairwise identification is calculated as the accuracy with which the predicted features (either the output of a translator or features extracted from the output of a generator) can correctly identify the corresponding true ones, in pairs consisting of a true sample and one of the remaining samples in the test set. We refer to the latter remaining sample as a candidate sample in the following. If the candidate sample belongs to the same category as the true sample, the identification is expected to be difficult. Conversely, if the candidate sample belongs to a different category from the true sample, identification becomes easier. This characteristic makes the metric highly dependent on categorical distinctions rather than the actual quality of feature prediction or reconstruction.
@@
Fig. 10: Expected pairwise sample identification performance in categorically structured data. The $x$-axis represents the number of categories in the test set. The $y$-axis represents the pairwise sample identification accuracy. Different lines represent various levels of accuracy ($q$) in distinguishing samples from different categories. The chance level of $0.5$ is represented by the bottom line. Above the graph, hypothetical similarity matrices are shown to illustrate the categorical structures of the samples, where the color (white/black) indicates similarity between samples. The block diagonal structure reflects the categorical nature of the data. Samples within the same category are assumed to be indistinguishable, resulting in a pairwise identification accuracy of $0.5$ (chance level). Samples from different categories can be distinguished with a pairwise identification accuracy of $q$, where q varies between $0.5$ and $1$.
@@
Here, we assume the test set comprises $k$ categories with test samples equally distributed across each category for simplicity. We model the situation mentioned above by setting the expected identification accuracy at the chance level (i.e., $0.5$) when a candidate sample belongs to the same category as the true one. Conversely, when the candidate sample belongs to a different category from the true sample, we set the expected identification accuracy to a parameter $q$, ranging from $0.5$ to $1$. This parameter $q$ reflects the ease of identification across categories. Assuming the number of test samples is sufficiently large, the pairwise identification accuracy $\mathrm{Acc}$ becomes
@@
(see Methods “Expected identification accuracy in imprecise reconstructions” for the derivation).
@@
Fig. 10 illustrates the relationship between pairwise identification accuracy and the number of categories in the test set through line plots of expected values. Inset figures depict the underlying similarity structure of the test set. Notably, even when identification within categories fails completely and succeeds only between two categories, pairwise identification accuracy can still reach a high value of up to $75$%. This highlights a major limitation: above-chance performance may simply reflect an ability to differentiate broad categories rather than accurately reconstruct crucial visual details. Indeed, the same pattern emerges even under a hold-out split when the samples in the held-out cluster can still be categorized broadly. Consequently, performance evaluation only relying on this metric can lead to misleading conclusions about the model’s actual capabilities and generalizability.
2.2.4 Preserved image information across hierarchical DNN layers
@@
The reconstruction of arbitrary visual images requires compositional latent features that can be effectively mapped into the image space. As our case study has suggested, hierarchical DNN features from VGG19 have been found to be suitable for zero-shot prediction or reconstruction tasks due to their compositional representations. At the same time, however, the extent to which these features truly map to the image space remains unclear. Indeed, a common narrative suggests that the hierarchical processing discards pixel-level information through progressively expanding receptive fields. Yet, this view is not entirely accurate.
@@
Fig. 11: Image recovery check for hierarchical latent features. The images show the ability to recover visual inputs from different layers of the VGG19 network. For each of the three sample images, three different optimization methods were applied. The left column shows results from pixel optimization, which directly optimizes pixel values to minimize the loss between the generated image features and the target image features (Mahendran and Vedaldi, 2015). The middle column displays optimization with a weak image prior, simultaneously optimizing both the weight parameters of the image prior models and their latent features (Ulyanov et al., 2018). The right column presents optimization with a pre-trained image prior, which optimizes the latent features of a parameter-fixed image generator model to minimize the loss between the output image features and the target image features (Dosovitskiy and Brox, 2016). Rows correspond to different layers of the VGG19 network, progressing from earlier layers at the top to later layers at the bottom. This progression illustrates how image information is preserved or lost at different stages of the network. Reasonable image recovery is possible even from relatively high-level layers of the VGG19 network, challenging the notion that higher layers in deep neural networks discard image-level visual information.
@@
For example, the latent features of auto-encoder models (Hinton and Salakhutdinov, 2006; Kingma and Welling, 2014; van den Oord et al., 2017) can represent images in a low-dimensional space while preserving their reversibility, which is reasonable considering that the model’s output is trained to match the input. Furthermore, Mahendran and Vedaldi (2015) showed that input images can be recovered with reasonable accuracy even from relatively high-level layers of a DNN designed for an object recognition task. It has also been argued in the neuroscience field that large receptive field sizes do not necessarily impair neural coding capacity as long as the number and density of units remain constant (Zhang and Sejnowski, 1999; Majima et al., 2017). These results challenge the notion that higher-level layers in DNNs discard all pixel-level information.
@@
To further illustrate this point, we performed a recovery check on each intermediate layer of the VGG19 network used in the iCNN methods (Fig. 11; see also Fig. 7). Given the DNN features of a target image, we optimized input pixel values to make the image’s latent features similar to the targets (see Methods “Recovery check of a single layer by iCNN”). We observed that input images can be recovered with reasonable accuracy from relatively high-level layers (around the 11th layer out of the total 19th layers). Furthermore, by introducing image generator networks to add constraints on image statistics, reasonable recovery can be achieved from even higher layers. By utilizing a weak image prior (Ulyanov et al., 2018), which contains only information about the structure of images without any prior information on natural images, input images can be recovered from the 13th layer. When using an image generator that has learned natural image information (Shen et al., 2019b; Dosovitskiy and Brox, 2016), input images can be recovered even from the 15th layer.
@@
Conversely, utilizing high-level image features, such as the output of DNNs, or features from other modalities, such as text annotations, is not a rational choice for visual image reconstruction tasks. These latent features make it challenging to recover the corresponding input image (Fig. 7B) and are insufficient as surrogate representations of perceived visual images. Although recent text-to-image models and predicted text features can easily generate images, the outputs should not be interpreted as reconstruction results. Instead, it is more appropriate to view them as visualizations of decoded semantic information. While such visualizations are valuable for illustrating purposes, it is crucial to recognize the significant distinction between semantic visualization and reconstruction.
2.2.5 How are we fooled by hallucinations of generative AIs?
Generative AIs have recently made remarkable progress, with models now capable of producing high-resolution and realistic images from text input (Ramesh et al., 2021) or generating text of a quality indistinguishable from human-written content (Brown et al., 2020). However, due to the complex internal structure of these models and the vast amounts of data they are trained on, we are often fooled by the outputs of generative AIs. For instance, when searching for an unfamiliar topic using a large language model (LLM) in our daily lives, we may not realize that the model is creating false concepts. This is likely a result of generative AIs being trained on a large amount of data and producing highly coherent and contextual responses. We may also believe these models are unbiased and can represent all possible data points, even though they inherently contain biases from their training data and developers (Messeri and Crockett, 2024). As we have observed, the generative AI-based reconstruction methods exhibit realistic appearances but poor generalizability (Fig. 2). Could similar issues occur in visual image reconstruction studies as well?
@@ 2.2.5 How are we fooled by hallucinations of generative AIs?
Traditionally, we have held two beliefs: (1) generating realistic images from brain activity is challenging, and (2) if the reconstruction pipeline effectively captures the brain’s representation under natural image perception, the model’s output should also appear realistic. Based on these beliefs, realistic reconstruction is often considered an indication of accurately reflecting the actual visual experience.
Fig. 12: The illustration of how realistic appearances can be misleading in visual reconstruction. On the left, Venn diagrams depict the relationship between two events: $R$, where the output has realistic appearances, and $T$, where the output truthfully reflects the visual experience. $T$ and a small dashed circle of $R$ depict a common pre-generative AI heuristic: the assumption that realistic appearance implies truthful reconstruction. Dashed arrows illustrate the effect of generative AI models (e.g., diffusion models) on this relationship. These models can potentially increase the overlap between $T$ and $R$, but may also expand areas of $R$ unrelated to $T$. On the right, a probabilistic interpretation of the relationship is presented. The equation represents the probability that the generator’s output truthfully reflects visual experience ($T$), given that the generator’s output has a realistic appearance ($R$). This conditional probability is derived from the areas in the Venn diagram. The expansion of $R$ without a proportional increase in $T$ can lead to a decrease in $\Pr[T\mid R]$, emphasizing the need for careful evaluation of reconstruction results beyond just assessing their realistic appearance.
@@
However, recent developments in generative AIs, such as diffusion models, have made it easy to produce convincing, realistic outputs, subverting the first assumption, i.e., $\Pr[R]\gg\Pr[\bar{R}]$. Consequently, it becomes invalid to infer that the visual images are accurately reflected in the generator outputs solely because they appear realistic. Rather, as shown in Fig. 12, the probability $\Pr[T|R]$ may become smaller as the generative AIs produce more convincing outputs (see also Fig. 7). This perspective emphasizes the need for careful evaluation of reconstruction performance, considering the possibility of hallucinations by generators. While pursuing realistic reconstructions to improve reconstruction fidelity is undoubtedly important, it would be counterproductive to obsess over naturalistic appearance to the point of neglecting the original goal of reconstructing perceived visual images.
3 Discussion
@@
While our study highlights the limitations of text-guided diffusion models for visual image reconstruction, it is important to acknowledge that these methods offer promising directions for brain decoding research. For instance, they can generate “ROI optimal stimuli,” which create images that activate a certain ROI activities maximally while not activating other ROIs activities, through the learned mapping between brain activities and latent features (Ozcelik et al., 2022; Ozcelik and VanRullen, 2023). Although current studies focus on well-known functional ROIs (e.g., face, word, place, and body regions), this approach can extend to less understood brain regions. By generating optimal stimuli and identifying robust visual patterns, we can formulate new hypotheses about functional representations in previously uncharacterized brain areas. This data-driven approach complements traditional hypothesis-driven methods, potentially uncovering novel functional regions overlooked by conventional analyses.
Moreover, while we emphasized the importance of zero-shot prediction, it is crucial to recognize that most brain decoding studies focus on classification tasks. Although these tasks are not zero-shot, they have nonetheless yielded valuable insights into neural representations (Haxby et al., 2001; Kamitani and Tong, 2005). In that sense, the text-guided or diffusion-based methods can also be utilized as tools for the visualization of decoded semantic contents (e.g., the supplementary movies in Horikawa et al., 2013). Such visualization can be highly useful for visually conveying decoded information, even if it does not constitute zero-shot prediction.
Additionally, it is worth noting that the individual components of text-guided reconstruction methods have already been utilized in various brain decoding applications. For instance, text latent features from deep neural networks and large language models (LLMs) have shown promise in analyzing semantic information from brain activity (Tang et al., 2023; Caucheteux et al., 2023; Zhou et al., 2024). Furthermore, diffusion models can extend beyond text-to-image generation as MindEye2 generates images from visual latent spaces. Notably, Cheng et al. (2023) successfully reconstructed subjective experiences using a diffusion model within a carefully designed experiment. Leveraging these components and exploring their potential synergies, researchers can advance brain decoding and visual image reconstruction while addressing the challenges highlighted in our study.
@@
The recent trend of collecting and sharing large-scale visual neural datasets, such as those by Hebart et al. (2023) and Xu et al. (2024), is a welcome development in the field of neuroscience. These datasets provide valuable resources for researchers to investigate brain function and advance our understanding of visual processing. The NSD is a particularly notable example, as it was created with the goal of extensively sampling brain responses to a wide range of natural visual stimuli (Allen et al., 2022; Naselaris et al., 2021). The NSD has been widely utilized in various studies (Prince et al., 2022; Gifford et al., 2023; Conwell et al., 2024), demonstrating its value to the research community. While our results suggest that the semantic and visual diversity of the NSD stimuli may not be as high as initially thought, and there is substantial overlap between the training and test sets provided by the NSD authors, this does not diminish the overall importance and usefulness of the dataset. However, to fully leverage the NSD and other publicly available large-scale datasets for developing generalizable and zero-shot prediction models, it is crucial to consider the data split between training and test sets carefully. While many large-scale datasets provide designated training and test splits, these splits are often not optimally designed to evaluate zero-shot prediction performance. When aiming for generalizable predictions beyond training examples as in visual image reconstruction, researchers should carefully verify whether significantly similar stimuli are included in both the training and test sets (Fig. A5). If significant overlap exists, redesigning the training-test split becomes necessary to ensure the test set contains stimuli substantially dissimilar from those in the training set, thereby enabling genuine evaluation of generalization capabilities (Fig. A7). Moreover, recent advancements in functional alignment and inter-site neural code conversion methods (Haxby et al., 2011; Yamada et al., 2015; Wang et al., 2024) hold promise for combining datasets from different sources, enabling truly larger-scale data analysis in neuroscience. These techniques allow researchers to align brain activity patterns across individuals and measurement sites even when stimuli are not shared across datasets. By leveraging these methods, researchers can pool data from various sources, increasing the sample size and diversity of the combined dataset, mitigating the limitations of individual datasets, and enhancing the development of generalizable and zero-shot prediction models.
@@
Investigating neural responses to natural stimuli is a highly valuable approach to understanding brain function and representation (Nastase et al., 2020; Hasson et al., 2020). As our brains have developed while being exposed to natural scenes, it is crucial to use natural stimuli, especially in model training. However, we should not forget that we are also capable of perceiving non-natural stimuli like artificial images. We would like to emphasize that there are potential pitfalls when relying too heavily on evaluations based solely on natural stimuli. With the increasing scale of neural data and the growing complexity of analysis pipelines, there is a risk that the learned mappings may produce unexpected shortcuts, just as we have demonstrated that the text-guided reconstruction methods exploited the semantic and visual overlap between training and test sets. In the field of comparative and developmental psychology, researchers often prioritize using not natural but simple stimuli for better experimental controls and more precise inferences about infants’ cognitive abilities (Kominsky et al., 2022; Frank, 2023). Drawing inspiration from this approach, we argue that the evaluation of visual image reconstruction should not be limited to complex natural stimuli alone. While natural stimuli are essential for ensuring ecological validity and understanding how the brain processes real-world information, it is equally important to assess the performance in controllable and transparent manners.
@@
We have observed that an inappropriate split between training and test stimulus sets can lead to spurious reconstruction, invalidating zero-shot predictions. It is crucial to recognize that these problems can arise in various research contexts. Several studies have used data where the stimuli shared the category information between the training and test sets, potentially compromising the validity of their results. Kavasidis et al. (2017) collected a dataset of EEG signals recorded during natural image perception. Their visual stimuli consisted of $2,000$ images selected from $40$ object categories in ImageNet ($50$ images per category) and the test set contained the same categories that are included in the training set (see also Li et al. (2018) and Xu et al. (2024) for other issues with the data set). Denk et al. (2023) attempted to develop music reconstruction methods from fMRI activity patterns. Their music stimuli consisted of $540$ music pieces selected from $10$ music genres, and the test set contained the same genres as the training set (Nakai et al., 2021). Orima et al. (2024) attempted to reconstruct perceived texture images from EEG signals. Their texture stimuli consisted of $191$ image patches extracted from $21$ natural textures, and they performed a reconstruction analysis in a leave-one-out manner. It should be carefully examined whether these studies may suffer from output dimension collapse, merely decoding the broad category-level information observed in the training set.
@@
Experimental design for training–test stimulus setups requires careful consideration. Dado et al. (2024) conducted an image reconstruction analysis from the multi-unit activity of a macaque using images generated from latent features of generative models. Importantly, all of the test stimuli were generated from the averaged latent features of the categories used in their training phase, suggesting the test stimuli are highly biased to the training set (see also Fig. A5). While the authors addressed potential biases by redesigning the training and test split, researchers should exercise prudence when utilizing the dataset. Our inspection of the movie stimuli from Nishimoto et al. (2011) revealed that many frames in the test movie stimuli were nearly identical to those in the training set (Fig. A11). This similarity likely results from temporally adjacent video frames being split between the training and test stimuli. While our preliminary analysis of their latent features (motion energy features) did not show unusual clustering (note also that the study presents a way of retrieving movie instances via a brain encoding model, rather than reconstruction in the current sense), caution is required when using the stimulus set to extract other types of features (Huth et al., 2012, 2016). Claims of predicting arbitrary instances or achieving zero-shot prediction warrant thorough scrutiny.
@@
There is a widespread practice of using test data for fine-tuning, which can also be a questionable procedure. Some studies have proposed methods that involve fine-tuning models using the entire test brain data, but not the test stimuli, following initial training with the training stimuli and brain data (Beliy et al., 2019; Chen et al., 2023). While the absence of test labels (stimuli) during fine-tuning may help avoid obvious overfitting, these approaches still treat test brain data as previously observed information, rather than as a proxy for novel data potentially encountered in real-world situations. Consequently, this practice violates test data independence, making it difficult to evaluate the model’s actual generalization ability. Although such procedures may effectively improve performance on existing benchmarks or competitions, it is crucial to recognize that incorporating test data information during any stage of training can undermine the validity of neuroscientific claims and limit the real-world applicability of the methods.
@@
The issue of double dipping, which refers to the use of the same dataset for both data/variable selection and selective analysis (inference and prediction), has been widely recognized in neuroscience (Kriegeskorte et al., 2009; Button, 2019). In classification tasks, while selecting input variables using test data is problematic, using the same output labels (target variables) for training and test sets is not inherently flawed, given that the nature of classification assumes consistent categories across datasets. However, the challenges we address in this study, although related to double dipping, present distinct concerns. In evaluating zero-shot prediction performance, the mere similarity of test labels (features) can lead to overestimating model performance. It is crucial to distinguish between conventional double dipping and the current issues we identify, such as output dimension collapse in zero-shot prediction scenarios. These emerging challenges necessitate careful consideration not only of input data independence but also of structural similarities between training and test sets in the output space.